# A Survey of Text Watermarking in the Era of Large Language Models

### v3

@@ 4.3. Watermarking during Token Sampling
The previous section primarily focused on incorporating watermarks during the logits generation phase for Large Language Models. However, even after embedding watermarks, it is necessary to sample the next token from the watermarked logits. Consequently, the effectiveness of the watermark might be influenced by different sampling methods; for instance, beam search typically allows for a higher watermark intensity. In this section, we will introduce a technique of watermarking during token sampling, which does not alter the logits produced by the LLM. The primary advantage of this method is that the generated text is usually unbiased, thus minimally affecting text quality and serving as the first line of defense against watermark forgery.
@@
To address these issues, Kuditipudi et al. (2023) proposed the use of a pseudo-random number sequence significantly longer than the text, randomly selecting a starting position from the sequence for each watermark insertion to introduce randomness. Additionally, during watermark detection, they incorporate a soft notion of edit distance (i.e., Levenshtein distance) into the computation of the alignment between text and the pseudo-random number sequence. This approach significantly enhances the robustness of the watermarking algorithm against watermark removal attacks. Apart from intervening in the sampling process of each token one by one, Hou et al. (2023) suggested incorporating watermarks during sentence-level sampling. The algorithm initially partitions the semantic embedding space into a watermarked region and a non-watermarked region, and then performs sentence-level rejection sampling until the sampled sentence falls within the watermarked region. Since the partition principles are based on sentence-level semantics, this approach significantly enhances robustness against watermark removal attacks such as paraphrasing. Currently, there is limited research on watermarking during token sampling, suggesting substantial potential for growth in this field. Moreover, the practical effectiveness and robustness of this method may require further validation through more experiments and real-world applications.
@@ 5. Evaluation Metrics for Text Watermarking
In sections 3 and 4, we have provided a comprehensive introduction to the existing text watermarking methods. Meanwhile, for a text watermarking algorithm, it is crucial to conduct a comprehensive evaluation from various perspectives. In this section, we will introduce how to evaluate a watermarking algorithm from four perspectives: success rate, text quality, robustness and unforgeability. Among these, success rate (section 5.1) refers to the ability to correctly detect watermarked texts. Text quality (section 5.2) assesses whether the quality of watermarked text is degraded compared to non-watermarked text. Robustness (section 5.3) examines whether the watermarked text can still be detected after watermark removal attacks. Finally, unforgeability (section 5.4) evaluates the difficulty for third parties to forge the watermark. Furthermore, Table 1 outlines the alignment between various text watermarking algorithms discussed in the survey and the evaluation perspectives they contribute to.
@@ 5.1.1. Zero-bit Watermark
In zero-bit watermarking, the watermarking algorithm can only determine whether a text contains a watermark, without the ability to extract additional information from the watermarked text. Given that the detection of a zero-bit watermark inherently represents a binary classification problem — discerning whether a text is watermarked or not — the evaluation of its success rate typically employs classification metrics, such as accuracy and F1 score. In most works (Kirchenbauer et al., 2023a; Zhao et al., 2023a; Liu et al., 2023b, c), since the proportion of watermarked and non-watermarked texts in the test dataset is usually equal, the values of accuracy and F1 score tend to be quite similar. Despite this, the F1 score is often more frequently utilized, accompanied by additional metrics including false positive and false negative rate (Kirchenbauer et al., 2023a; Zhao et al., 2023a; Liu et al., 2023b). Here, false positives denote the erroneous classification of non-watermarked texts as watermarked, whereas false negatives refer to the incorrect classification of watermarked texts as non-watermarked. Generally, false positives are more important since misidentifying human-generated texts as watermarked can lead to more adverse consequences.
However, since most watermark detection algorithms require a threshold to determine the F1 or accuracy values, the uncertainty of this threshold may introduce unfairness in the performance comparison of different algorithms. Consequently, some works (Zhao et al., 2023a; Liu et al., 2023c) have reported F1 values at fixed false positive rates of 1% and 10%, while others have reported the optimal F1 score across all potential thresholds to ensure a fairer comparison (Liu et al., 2023c). In addition, some works employ hypothesis testing methods to calculate the p-value as an important metric for sucess rate. Different algorithms utilize various hypotheses to compute the p-value. For example, Kirchenbauer et al. (2023a) involves determining if the calculated z-score exceeds a certain threshold, while (Kuditipudi et al., 2023) hypothesized that the key generating the watermark text has a higher probability of detecting the watermark compared to other randomly selected keys. This method (Kuditipudi et al., 2023) does not require a predefined threshold but necessitates multiple runs of the detection algorithm, which could slow down the detection speed.
@@ 5.1.2. Multi-bit Watermark
For multi-bit watermarking methods (Wang et al., 2023b; Rizzo et al., 2016; Abdelnabi and Fritz, 2021; Yang et al., 2022; Yoo et al., 2023b, b, a), they not only detect the presence of a watermark but also extract more detailed watermark information. For instance, a watermarked text might convey specific data like This text is generated by GPT-4 on June 6 by the Administrator (Wang et al., 2023b). When evaluating the success rate of multi-bit watermarks, it is crucial not only to assess their accuracy in extracting watermark information but also to consider the payload of the watermark, which refers to the number of bits in the watermark information.
Assume that a piece of watermark information, $w$, can be represented using $n$ bits, denoted as $w=b_{1}b_{2}…b_{n}$, where each $b_{i}$ can take a value of 0 or 1. In this context, the most commonly used metric for assessing success rate is the bit error rate (BER) (Yoo et al., 2023b), which is the probability of incorrectly predicting each bit during detection, or the bit accuracy (Yoo et al., 2023a; Abdelnabi and Fritz, 2021), which is the rate of correctly predicting each bit. These are two complementary metrics. Typically, the calculation of the bit error rate is conducted under a fixed bit number, and as the bit number increases, the bit error rate also tends to increase, eventually approaching 50% (random).
Therefore, the bit capacity of a watermark algorithm has become an important evaluation criterion, commonly referred to as payload (Yang et al., 2022), Bits Per Watermark (BPW) (Yang et al., 2022; Yoo et al., 2023b), or code rate (Wang et al., 2023b; Rizzo et al., 2016). The payload can be calculated by dividing the total number of bits in the watermark information by the number of tokens. For a watermark algorithm, there is an upper limit to its payload, and enhancing the payload typically comes at the cost of either reduced text quality (section 5.2) or decreased robustness (section 5.3). Additionally, the value of the payload is contingent on the textual context; in scenarios with higher entropy, the payload tends to be higher, whereas in lower entropy scenarios, the payload is usually lower.
@@ 5.2.1. Perplexity
The perplexity of watermarked text is generally higher than that of non-watermarked text, indicating a slight decrease in text quality. Generally, the impact of watermarking algorithms on text quality is correlated with the strength of the watermark. The higher the watermark strength, the more evident the decline in text quality. For instance, in the KGW algorithm (Kirchenbauer et al., 2023a), a higher value of $\delta$ results in greater impact on text quality. Takezawa et al. (2023) suggest that for longer texts, a weaker watermark strength can be employed to minimize the effect on text quality while maintaining the effectiveness of the watermark.
@@ 5.2.2. Semantic Score
Although text perplexity facilitates the evaluation of textual consistency and fluency, it could not assess the accuracy of watermarked texts, specifically in terms of semantic consistency between watermarked and un-watermarked text. Consequently, some studies employ semantic scores, which measure the semantic similarity between watermarked and non-watermarked texts, to evaluate the impact of watermarking algorithms on text quality.
@@
Although semantic scores assessment based on Natural Language Inference (NLI) offers an in-depth semantic analysis, it might still fall short in accurately capturing variations at the level of individual words or phrases. To address this, Zhang et al. (2023a) employed BERT-Score (Zhang et al., 2019) for a word-level detailed comparison of texts. BERT-Score is more adept at evaluating whether the watermark has altered specific vocabulary or expressions in the original text.
@@ 5.2.3. Task-specified Evaluation
Although the assessment method based on semantic scores could effectively evaluate whether adding watermarks alters the text semantics, its impact on real-world applications remains unclear. Consequently, many studies are now focusing on exploring the effects of watermarking algorithms on specific downstream tasks to assess their impact on text quality. These specific downstream tasks include machine translation (Topkara et al., 2006a; Hu et al., 2023; Wu et al., 2023; Zhao et al., 2023b), sentiment classification (Yang et al., 2023), knowledge understanding (Tu et al., 2023), code generation (Lee et al., 2023; Sun et al., 2023; Tu et al., 2023), text summarization (Hu et al., 2023; Wu et al., 2023; Tu et al., 2023), story generation (Zhao et al., 2023b), question answering (Tu et al., 2023) and instruction following (Tu et al., 2023), as illustrated in Figure 6.
Machine Translation. Typically, only watermarking Large Language Models (Section 4) incorporate machine translation as a downstream task for testing text quality. Specifically, the evaluation involves comparing the translation outcomes between a watermarked LLM and the original unwatermarked LLM. This comparison is conducted using the BLEU score, a widely used metric in machine translation. For the choice of translation LLMs, (Hu et al., 2023; Wu et al., 2023) employed the Multilingual BART (Liu et al., 2020), while Takezawa et al. (2023) utilized the NLLB-200 model (Costa-jussà et al., 2022). The translation data typically employed is the WMT14 dataset, where the translations between French, German, and English are the most commonly utilized settings. Most watermarking approaches for LLMs result in a slight decrease in BLEU scores (Kirchenbauer et al., 2023a; Takezawa et al., 2023; Liu et al., 2023b). However, the unbiased watermark methods (Hu et al., 2023; Wu et al., 2023) exhibit almost no decline in BLEU values, demonstrating the superiority of unbiased watermarking.
@@
Code Generation. Text watermarking for code generation is an important application, which could test the impact of watermarking on code functionality. Code evaluation could use unit test metrics like pass@k (Kulal et al., 2019), or matching metrics such as BLEU and Exact match. Sun et al. (2023) inserted watermark features into a subset of a dataset to achieve code dataset watermarking. They showed negligible differences in BLEU and exact match scores between models trained on datasets with and without watermarks. However, embedding watermarks into individual code is more challenging than watermarking an entire dataset. Lee et al. (2023) added watermarks to large models for code generation, leading to a nearly 10% performance decline in pass@100 and pass@80 metrics. Given that even minor modifications can disrupt code functionality, and considering code’s inherently low entropy, the code generation task presents a highly challenging downstream task for text quality test (Tu et al., 2023).
Text Summarization. Similar to machine translation scenarios, only the watermark algorithms for LLM consider text summarization as a downstream evaluation task. Specifically, it compares the effectiveness of text summarization between a watermarked Large Language Model (LLM) and an unwatermarked LLM. The common evaluation metric for text summarization is ROUGE (Lin, 2004). Furthermore, the most frequently used large model for summarization is BART-Large (Liu et al., 2020), with the CNN-DM dataset (Hermann et al., 2015) being prevalent. In practical results, current watermark algorithms have a relatively minimal impact on text summarization tasks. The algorithm by Kirchenbauer et al. (2023a). only causes a slight decrease in ROUGE scores, whereas the unbiased watermark (Hu et al., 2023; Wu et al., 2023)algorithm hardly affects the ROUGE scores.
Story Generation. Similar to text summarization tasks, story generation also presents a suitable scenario for evaluating watermark algorithms for Large Language Models. Tests in the story generation context typically involve inputting the first half of a story and having the model (LLM) predict its ending. The ROCstories dataset (Mostafazadeh et al., 2016) is commonly used, with ROUGE as the evaluation metric. According to the experiments by Zhao et al. (2023b), current watermark algorithms still cause a 1%-2% decrease in performance on ROUGE scores. Furthermore, minimal research has been conducted on story generation performance, indicating potential for future exploration.
Question Answering. Question answering (QA) is an important downstream application of Large Language Models, and testing watermark algorithms for LLMs on this task is equally crucial. Tu et al. (2023) conducted tests on three different QA tasks, specifically: the ELI5 dataset (Fan et al., 2019) for long-form QA tasks, the FinQA dataset (Maia et al., 2018) for Finance QA tasks, and the HotpotQA dataset (Yang et al., 2018) for multi-hop reasoning QA tasks. For the long-form QA and Finance QA tasks, the Rouge-L metric was used for evaluation, while the F1 score was utilized for the multi-hop reasoning QA task. Experimental results revealed that, after the introduction of watermarks, all current watermark algorithms experienced a performance decline of about 50% in Question Answering tasks, indicating the challenging nature of watermark algorithms in the context of Question Answering.
Instruction Following. The instruction following ability of Large Language Models has recently become an especially important aspect to assess, reflecting whether LLMs can accurately follow user instructions to generate output in open-ended situations. Tu et al. (2023) tested the impact of the current watermark for LLMs algorithm on the Instruction Following task using the AlpacaFarm dataset (Dubois et al., 2023). The evaluation method adopted was GPT4-Judge (Zheng et al., 2023), where the GPT-4 model judges which output, between the watermarked LLM and Davinci-003, is better in response to a given instruction. Under this metric, both Llama2-7B-chat and Internlm-7B-8k models showed over a 90% decline in performance with watermark, indicating that instruction following presents a particularly challenging scenario for watermark algorithms.
@@
5.2.4. Output Diversity.
Although previous text quality assessment methods provide comprehensive evaluations of consistency, fluency, and accuracy, they still overlook an essential aspect: assessing the diversity of watermarked texts. Diversity evaluation is often targeted at the watermark algorithms for Large Language Models, since these algorithms involve embedding watermarks into LLMs, necessitating an assessment of whether the diversity of the text generated by the watermarked LLMs has changed. Text diversity is defined by calculating the proportion of unique n-grams in a text sequence, with the formula being the negative logarithm multiplied by the product of (1 - proportion) of unique n-grams from 1 to N. A higher diversity score indicates fewer repeated n-grams and richer text. The specific formula is defined as follows:
@@
In the work Kirchenbauer et al. (2023b), it was discovered that for the KGW watermarking algorithm (Kirchenbauer et al., 2023a), the context width (i.e., the number of tokens used on the left to hash and generate the green list) has the most significant impact on text diversity. A larger context width enhances the diversity of the text, but this comes at the cost of reduced robustness of the watermark to text modifications. Furthermore, with a larger context width, as the watermark strength increases, so does the diversity. Conversely, with a smaller context width, an increase in watermark strength leads to a decrease in diversity. Currently, only a few studies on watermarking for Large Language Models have assessed its impact on text diversity. We suggest that future work could focus more on evaluating the effect of watermarking on diversity.

### v4

@@ 4.3. Watermarking during Token Sampling
The previous section primarily focused on incorporating watermarks during the logits generation phase for Large Language Models. In this section, we will introduce a technique of watermarking during token sampling, which does not alter the logits but utilize watermark message to guide the sampling of each token. The main advantage of this method is unbiased text generation, which minimally affects text quality and offers an initial defense against watermark forgery.
@@
To address these issues, Kuditipudi et al. (2023) proposed the use of a pseudo-random number sequence significantly longer than the text, randomly selecting a starting position from the sequence for each watermark insertion to introduce randomness. Additionally, during watermark detection, they incorporate a soft notion of edit distance (i.e., Levenshtein distance) into the computation of the alignment between text and the pseudo-random number sequence. This approach significantly enhances the robustness of the watermarking algorithm against watermark removal attacks. Apart from intervening in the sampling process of each token one by one, Hou et al. (2023) suggested incorporating watermarks during sentence-level sampling. The algorithm initially partitions the semantic embedding space into a watermarked region and a non-watermarked region, and then performs sentence-level rejection sampling until the sampled sentence falls within the watermarked region. Since the partition principles are based on sentence-level semantics, this approach significantly enhances robustness against watermark removal attacks such as paraphrasing. Current research in token-sampling watermarking is limited, indicating room for advancement. The effectiveness and robustness of these methods warrant further exploration through experiments and real-world applications.
@@ 5. Evaluation Metrics for Text Watermarking
In Sections 3 and 4, we presented a detailed overview of existing text watermarking techniques. A thorough evaluation of a text watermarking algorithm is essential. This section outlines the evaluation metrics from multi perspectives: success rate, text quality, robustness, and unforgeability. Among these, success rate (Section 5.1) measures the accuracy of detecting watermarked texts. Text quality (Section 5.2) evaluates the integrity of watermarked text in comparison to its original form. Robustness (Section 5.3) assesses the detectability of watermarks after removal attacks. Unforgeability (Section 5.4) considers the difficulty in counterfeiting the watermark. Additionally, Table 1 outlines the alignment between various text watermarking algorithms discussed in the survey and the evaluation perspectives they contribute to.
@@ 5.1.1. Zero-bit Watermark
In zero-bit watermarking, the algorithm discerns the presence of a watermark in text but cannot extract additional information. This detection essentially constitutes a binary classification problem, evaluating whether text is watermarked. Classification metrics like accuracy and F1 score are typically used for evaluation. In many studies (Kirchenbauer et al., 2023a; Zhao et al., 2023a; Liu et al., 2023b, c), given the equal distribution of watermarked and non-watermarked texts in test datasets, accuracy and F1 score values are often comparable. However, F1 score, along with false positive and false negative rates, is more frequently employed (Kirchenbauer et al., 2023a; Zhao et al., 2023a; Liu et al., 2023b). False positives signify misclassifying non-watermarked texts as watermarked, and false negatives involve incorrectly identifying watermarked texts as non-watermarked. False positives are critical due to misidentifying human-generated texts as watermarked can lead to more adverse consequences.
Nevertheless, the reliance on a threshold for determining F1 or accuracy introduces comparability challenges among algorithms. Some studies (Zhao et al., 2023a; Liu et al., 2023c) report F1 scores at fixed false positive rates of 1% and 10%, while others (Liu et al., 2023c) present optimal F1 scores across all thresholds for a fair comparison. Additionally, hypothesis testing methods are used to calculate the p-value as a key success rate metric. Different algorithms employ varying hypotheses for p-value computation. For instance, Kirchenbauer et al. (2023a) assess if the calculated z-score surpasses a threshold, whereas Kuditipudi et al. (2023) hypothesizes that the key generating the watermark text detects the watermark with higher probability than random keys. This approach, not requiring a predefined threshold, needs multiple detection runs, potentially slowing detection speed.
@@ 5.1.2. Multi-bit Watermark
In multi-bit watermarking methods (Wang et al., 2023b; Rizzo et al., 2016; Abdelnabi and Fritz, 2021; Yang et al., 2022; Yoo et al., 2023b, a), the focus extends beyond merely detecting watermarks to extracting detailed watermark message. For example, a watermarked text may encode specific data, such as "This text is generated by GPT-4 on June 6 by the Administrator" (Wang et al., 2023b). Evaluating the efficacy of multi-bit watermarks involves not only assessing accuracy in information retrieval but also considering the watermark’s payload, the number of bits in the watermark message.
Consider watermark message $w$ encoded in $n$ bits, represented as $w=b_{1}b_{2}…b_{n}$, where each $b_{i}$ is binary. The prevalent metrics for success rate evaluation are the bit error rate (BER) (Yoo et al., 2023b), indicating the likelihood of erroneous bit predictions, and bit accuracy (Yoo et al., 2023a; Abdelnabi and Fritz, 2021), the rate of correct bit predictions. These metrics are complementary. BER is usually calculated for a fixed bit number; as the number of bits increases, BER tends to rise, potentially reaching 50% (random).
Therefore, the bit capacity, or payload, of a watermark algorithm is a critical evaluation metric, known as Bits Per Watermark (Yang et al., 2022; Yoo et al., 2023b) or code rate (Wang et al., 2023b; Rizzo et al., 2016). Payload is calculated by dividing the total bit count of the watermark information by the number of tokens. The payload of a watermark algorithm has an upper limit, with enhancements often compromising text quality (section 5.2) or robustness (section 5.3). Furthermore, the payload is context-dependent: higher entropy scenarios yield greater payloads, whereas lower entropy scenarios result in smaller payloads.
@@ 5.2.1. Perplexity
Watermarked texts typically exhibit higher PPL, reflecting a quality reduction. The higher the watermark strength, the more evident the decline in text quality. For example, in the KGW algorithm (Kirchenbauer et al., 2023a), an increased $\delta$ value noticeably affects text quality. Takezawa et al. (2023) recommend using lower watermark strength for longer texts to mitigate quality degradation while preserving watermark efficacy.
@@ 5.2.2. Semantic Score
Although text perplexity facilitates the evaluation of textual consistency and fluency, it could not assess the accuracy of watermarked texts, specifically in terms of semantic consistency between watermarked and non-watermarked text. Consequently, some studies employ semantic scores, which measure the semantic similarity between watermarked and non-watermarked texts, to evaluate the impact of watermarking algorithms on text quality.
@@
Although semantic scores assessment based on Natural Language Inference (NLI) offers an in-depth semantic analysis, it might still struggle to accurately capture variations at the level of individual words or phrases. To address this, Zhang et al. (2023a) employed BERT-Score (Zhang et al., 2019) for a word-level detailed comparison of texts. BERT-Score is more adept at evaluating whether the watermark has altered specific vocabulary or expressions in the original text.
@@ 5.2.3. Task-specified Evaluation
Although the assessment method based on semantic scores could effectively evaluate whether adding watermarks alters the text semantics, its impact on real-world applications remains unclear. Consequently, many studies are now focusing on exploring the effects of watermarking algorithms on specific downstream tasks to assess their impact on text quality. These tasks include machine translation (Topkara et al., 2006a; Hu et al., 2023; Wu et al., 2023; Zhao et al., 2023b), sentiment classification (Yang et al., 2023), knowledge understanding (Tu et al., 2023), code generation (Lee et al., 2023; Sun et al., 2023; Tu et al., 2023), text summarization (Hu et al., 2023; Wu et al., 2023; Tu et al., 2023), story generation (Zhao et al., 2023b), question answering (Tu et al., 2023), and instruction following (Tu et al., 2023), as shown in Figure 6.
Machine Translation. Commonly, the assessment of watermarking in Large Language Models (Section 4) includes machine translation as a downstream task to evaluate text quality. This evaluation involves comparing the translation outputs from watermarked and original non-watermarked LLMs, utilizing the BLEU score, a standard metric in machine translation. For translation LLMs, (Hu et al., 2023; Wu et al., 2023) utilized Multilingual BART (Liu et al., 2020), and Takezawa et al. (2023) employed the NLLB-200 model (Costa-jussà et al., 2022). The WMT14 dataset, particularly translations between French, German, and English, is predominantly used for testing. Watermarking LLMs typically leads to a marginal reduction in BLEU scores (Kirchenbauer et al., 2023a; Takezawa et al., 2023; Liu et al., 2023b). Nonetheless, unbiased watermark methods (Hu et al., 2023; Wu et al., 2023) show negligible impact on BLEU scores, indicating their efficacy in maintaining translation quality.
@@
Code Generation. Text watermarking for code generation is an important application, which could test the impact of watermarking on code functionality. Code evaluation could use unit test metrics like pass@k (Kulal et al., 2019), or matching metrics such as BLEU and Exact match. Sun et al. (2023) implemented watermarking in a code dataset by embedding features in a dataset subset. They reported minimal differences in BLEU and Exact match scores between models trained on datasets with and without watermarks. Nonetheless, watermarking individual code pieces is more complex than watermarking an entire dataset. Lee et al. (2023) incorporated watermarks into large models for code generation, resulting in approximately a 10% decrease in pass@100 and pass@80 metrics. Considering the sensitivity of code to minor alterations, preserving the code utility while achieving watermarking poses a significant challenge(Tu et al., 2023).
Text Summarization. Only the watermark algorithms for LLM consider text summarization as a downstream evaluation task. Specifically, it compares the effectiveness of text summarization between a watermarked Large Language Model (LLM) and an non-watermarked LLM. The evaluation metric for text summarization is ROUGE (Lin, 2004). Furthermore, the most frequently used large model for summarization is BART-Large (Liu et al., 2020), with the CNN-DM dataset (Hermann et al., 2015) being prevalent. In practical results, current watermark algorithms have a relatively minimal impact on text summarization tasks. The algorithm by Kirchenbauer et al. (2023a) only causes a slight decrease in ROUGE scores, whereas the unbiased watermark (Hu et al., 2023; Wu et al., 2023)algorithms hardly affect the ROUGE scores.
Story Generation. Similar to text summarization tasks, story generation also presents a suitable scenario for evaluating watermark algorithms for Large Language Models. Tests in the story generation context typically involve inputting the first half of a story and having the model (LLM) predict its ending. The ROCstories dataset (Mostafazadeh et al., 2016) is commonly used, with ROUGE as the evaluation metric. According to the experiments by Zhao et al. (2023b), current watermark algorithms still cause a 1%-2% decrease in performance on ROUGE scores.
Question Answering. Question answering (QA) is an important downstream application of Large Language Models. Tu et al. (2023) conducted tests on three different QA tasks, specifically: the ELI5 dataset (Fan et al., 2019) for long-form QA tasks, the FinQA dataset (Maia et al., 2018) for Finance QA tasks, and the HotpotQA dataset (Yang et al., 2018) for multi-hop reasoning QA tasks. For the long-form QA and Finance QA tasks, the Rouge-L metric was used for evaluation, while the F1 score was utilized for the multi-hop reasoning QA task. Experimental results revealed that, after the introduction of watermarks, all current watermark algorithms experienced a performance decline of about 50% in QA tasks, indicating the challenging nature of watermark algorithms in the context of QA.
Instruction Following. The capacity of Large Language Models (LLMs) to adhere to user instructions in generating responses for open-ended tasks is increasingly crucial. Tu et al. (2023) examined the influence of the prevalent LLM watermarking algorithm on Instruction Following using the AlpacaFarm dataset (Dubois et al., 2023). The assessment utilized GPT4-Judge (Zheng et al., 2023), wherein GPT-4 evaluates the appropriateness of outputs from the watermarked LLM and Davinci-003 in response to specific instructions. According to this criterion, both the Llama2-7B-chat and Internlm-7B-8k models exhibited over a 90% decrease in performance when watermarked, highlighting the significant challenge watermark algorithms face in instruction following tasks.
5.2.4. Text Diversity.
Although previous text quality assessment methods provide comprehensive evaluations of consistency, fluency, and accuracy, they still overlook an essential aspect: assessing the diversity of watermarked texts. Diversity evaluation is often targeted at the watermark algorithms for Large Language Models. These algorithms embed watermarks into LLMs, necessitating an assessment of evaluating whether the diversity of the text generated by LLMs has changed. Text diversity is defined by calculating the proportion of unique n-grams in a text sequence, with the formula being the negative logarithm multiplied by the product of (1 - proportion) of unique n-grams from 1 to N. A higher diversity score indicates fewer repeated n-grams and richer text. The specific formula is defined as follows:
@@
In Kirchenbauer et al. (2023b), it was noted that for the KGW algorithm (Kirchenbauer et al., 2023a), context width—the number of tokens used to hash and create the green list—significantly influences text diversity. A larger context width enhances the diversity of the text, but this comes at the cost of reduced robustness of the watermark to text modifications. Furthermore, with a larger context width, as the watermark strength increases, so does the diversity. Conversely, with a smaller context width, an increase in watermark strength leads to a decrease in diversity. Presently, limited research on LLM watermarking addresses its impact on text diversity. Future studies should focus more on this aspect of watermarking evaluation.